# Data Requirement Goal Modeling for Machine Learning Systems

### v1

@@ 3.3. Customization mechanism of the DRGM
The customization mechanism is designed to adapt the initial DRGM to different ML problems and contexts. In addition, it assists requirements engineers, particularly those who are non-experts in ML, in selecting between datasets or evaluating a dataset to ensure it satisfies the data requirements for a given ML problem, as illustrated in Fig. 4
Figure 4. The customization mechanism.
@@
Given that data requirements are highly dependent on the type of ML problem and are often context-specific, the customization mechanism categorizes data requirements goals into three sets. The first set includes goals that are of high importance across all ML problems, referred to as goal set #1. The second set comprises goals, subgoals, and tasks that vary based on the ML type (e.g., Regression, Classification, Time Series). The third set pertains to data requirements that vary based on the collected data and the specific context. The contents of each set are detailed in Table 2.
Table 2. The customization mechanism sets
@@
3.3.1. Goals with Fixed Importance
@@
Examining the factors that determine set #1, Data Quantity is consistently assigned a ”High” importance, as having more high-quality data generally leads to better results (r2, ). To satisfy the Data Quantity goal, both Data Availability and Data Accessibility must also be satisfied. Data Quality is equally critical; if Data Quality is not met, the acquired data will be ineffective, resulting in a ”garbage-in, garbage-out” scenario, as noted by a data scientist in the study (r2, ). Similarly, Data Representativeness, Data Completeness, and Data Accuracy hold high importance across all contexts. Although Data Consistency and Data Safety are also important, their significance is comparatively lower, and their importance is therefore set to ”Medium.”
3.3.2. Goals with Different Importance Based on ML Problem Type
@@
Two elements related to data requirements vary in importance depending on the type of ML problem: Data Balancedness and Data Availability. Data Balancedness is highly important for classification problems but holds moderate importance for regression problems. For classification problems, techniques such as SMOTE, oversampling, and undersampling can be applied to address imbalanced data (r15, ; r14, ). For regression problems, data transformations to handle imbalance, often referred to as skewness, can be used (r16, ; r17, ). To set KPIs for Data Balancedness in classification problems, users should evaluate the balance as a percentage from the ration (e.g., for binary classification, a 1:1 class ratio corresponds to 100% balance). For regression problems, the Shapiro-Wilks test is suggested. This test’s null hypothesis states that the data is drawn from a normally distributed population. If $P>0.05$, the null hypothesis is accepted, and the data is considered normally distributed. If the null hypothesis is rejected, the target variable is highly skewed (r18, ). The initial values for these KPIs are provided in Table 3, and evaluation values are determined based on the percentages or test results.
Table 3. Balancedness KPI values
ML problem Worst Value Threshold Value Target Value Unit
Classification 0 50 100 percentage
Regression 0.05 0.05 0.05 p-test score
@@
For setting KPIs for Data Availability, several approaches exist to calculate the required amount of data based on the learning rate (r19, ). However, these approaches require an initial dataset and a pre-built ML algorithm. As such, they are primarily useful for determining the additional data needed during a pilot study for a selected dataset. When no initial dataset exists, there are no solid, theoretically proven rules for determining how much data an ML algorithm requires for training. Nonetheless, a heuristic known as the ”rule of 10,” based on practitioners’ experiences, is frequently mentioned in the gray literature (r21, ; r22, ; r20, ) and in a white paper for classification problems (r23, ). This rule suggests using ten instances per class or ten instances per predictor. Variations of this rule propose reducing the requirement to five instances or increasing it to 100. For this work, the KPI is set following this rule, incorporating its three variations.
In the case of image classification using deep learning, if a pre-trained model is not used, 1000 images per class are typically required as the minimum threshold (r24, ). However, leveraging pre-trained models significantly reduces this requirement, with as few as 20 images per class being sufficient in the worst-case scenario. Additional KPI values are derived from practitioners’ experiences and estimates.
@@
For forecasting with time-series data, particularly in seasonal data (e.g., weather or sales data), an ideal guideline suggests using the number of seasons within a year plus five additional data points (r25, ). However, real-world scenarios often involve randomness, requiring the model to utilize multiple ”seasons” worth of data. Based on (9718423, ) and expert input, the worst-case requirement is set to one times the number of data points in a season, the threshold to two times, and the maximum to ten times the number of data points in a season. For univariate, non-seasonal data using statistical models, the latest 50 observations are generally sufficient for short-term forecasts (r27, ; r28, ). Some practitioners suggest $40$ observations may suffice, while others recommend using up to $100$ for more accurate results. For multivariate time-series predictions, the ”rule of 10” is also applied to refine the estimate.
A detailed explanation of how to set the data size KPI for various ML problem types is provided in Table 4, and the process is summarized in the MLProblemTypeBased UML activity diagram shown in Fig. 5.
Table 4. KPI values for data size
@@
Classification – Image – pre-trained model 20 x #classes 100 * #classes 1000 * #classes Number of images
Classification – Image – starting from scratch 500 * #classes 1000 * #classes 10,000 * #classes Number of images
@@
Time Series - Seasonal 1 2 10
Years (or the maximum
seasonality period) worth of data
@@
Figure 5. ML problem type based UML Activity diagram
3.3.3. Goals with Different Importance Related to Context
Many data requirements are context-dependent. For the Data Representativeness goal, spatial data or seasonal-temporal data must be well-represented. If the prediction outcome directly impacts human subjects, it is critical to ensure Data Representativeness across the targeted demographics to avoid discrimination.
@@ 3.3.3. Goals with Different Importance Related to Context
Regarding Data Management, its importance is set to ”Low” only if the model is built once and never updated. In such cases, Data Freshness is also set to ”Low,” and data does not need to be sourced from sustainable resources. However, a study by (r32, ) found that one-third of models require updates at least monthly, and nearly one-quarter require daily tuning. This is especially relevant in fields such as marketing, stock market forecasting, and short-term weather prediction. For models requiring regular updates, Data Management is assigned ”High” importance. Similarly, Data Freshness is set to ”High” for models requiring frequent tuning, and the data must be acquired from sustainable sources. For models with irregular updates, the importance of Data Freshness and sourcing sustainable data is set to ”Medium.”
@@
When considering Data Ethics, the content of the dataset must first be evaluated. If the data identifies human subjects, identifying information must be removed or protected to prevent discrimination. Additionally, user consent must be obtained as a task under Data Privacy. GDPR regulations impose stricter consent requirements for identifying information, such as a photograph of a person’s house. These rules apply when the subjects are European Union (EU) citizens or if the system will be used within the EU (r29, ). In other regions, compliance should align with the relevant local regulations.
@@
Next, we examine the Data Privacy goal. If the dataset contains sensitive information, such as health records, financial records, university records, or business records, the importance of both Data Privacy and Data Security is set to ”High.” If the data is private but not sensitive, their importance is set to ”Medium.” For public data, the importance of Data Privacy and Data Security is set to ”None.” For private data, a release form must be signed to obtain access. For medical data, approval from an International Review Board (IRB) is required. Additionally, the importance of the goal Data Free from Discrimination is set to ”High” if the ML model’s decisions have a significant impact on human lives, such as in job recruitment (r30, ) or prison release (r31, ). In such cases, data should be obtained from authoritative sources, and sensitive demographic fields should either be removed or designated as protected to prevent bias in the system.
Finally, the overall importance of Data Ethics is determined by the highest importance level among its subgoals. Before initiating the evaluation, an analyst should consult with both domain and legal experts to address any potential misrepresentation in the data and any legal concerns specific to the field. A summary of this process is illustrated in the ContextBased UML activity diagram, Fig. 6.
Figure 6. Context Based Activity Diagram
3.3.4. Dataset evaluation and selection
@@
After customizing the DRGM based on the problem type and context, the dataset(s) are evaluated using the GRL evaluation strategy to select the dataset that best satisfies the Data Actor. We propose an evaluation scale ranging from 0 to 100. During the evaluation, all unrelated leaf elements should have their qualitative evaluation values set to ”satisfied,” and their negative contribution links removed to avoid negatively impacting the overall evaluation.
@@
4. Illustrative Examples for Using DRGM
@@
This section shows the customization of the DRGM approach using the proposed customization mechanism. Since the data requirements depend on the ML prediction type and are mostly context-related, we choose two different illustrative examples. The first one in the medical field use classification as a type of ML. The second example uses the regression for time series weather forecasting.
Figure 7. Anemia Detection Customized DRGM
Figure 8. Anemia Detection Dataset#1 Evaluation Model
Figure 9. Anemia Detection Dataset#2 Evaluation Model
4.1. Non-invasive Diagnosis of Anemia System using Machine Learning
@@
Context. Researchers in (r33, ) aimed to build an ML system to diagnose anemia using a video of a patient’s fingertip. To collect the training and testing dataset, three infrared LEDs and one white LED were positioned around the camera capturing the video. The LEDs were activated sequentially, and the average values of each of the three channels (RGB) from the video were calculated separately. These values were then used to train an ML model with the target classes being severely anemic, moderately anemic, mildly anemic, or healthy in a multi-class classification problem.
@@
Customizing the DRGM. First, we follow the ML problem type based UML activity diagram and as we are dealing with a classification problem we set the balancedness to ”HIGH”, added the KPI as per the diagram, then added the related task. As for the KPIs for the data availability, we have $4$ classes, so we set the worst value to $20$, threshold value to $40$, and target value to $400$. We then follow the context-based UML activity diagram, and as the data relates to human subject Data Ethics requirements will be set to HIGH. No special legal requirements other than the IRB and the consent release forms are required. Different age groups, both female and male, and different skin colors should be represented. By completing this step, we now have a customized DRGM based on the ML type and based on the context, Fig. 8.
Dataset Evaluation. The initial plan for the dataset involved collecting data in-house from 100 participants. Due to time constraints, the goal was to create a representative dataset and address any imbalance through sampling. Redundant data would be removed, and since the model was intended to be built only once, data would not be sourced from sustainable resources. Evaluating the planned dataset using the customized DRGM, shown in Fig. 8, indicated promising outcomes if the implementation were executed properly.
@@
However, as the implementation of the project described in (r33, ) progressed, the total number of participants was reduced to $80$. The conventional participant selection process led to a dataset that lacked adequate representation of individuals with fair skin tones, male participants, and those who were moderately or severely anemic. After performing undersampling on the healthy class, the dataset remained highly imbalanced, with class ratios of 3:5:12:40. Consequently, the balancedness KPI was set to $20$. Certain inconsistencies in the dataset could not be resolved, as recordings were conducted in different rooms. Additionally, some accuracy issues arose due to the fading of lights towards the end of the data collection phase. The presence of multiple participants in the recording rooms further posed risks to data safety. The evaluation of the collected dataset, presented in Fig. 9, highlights why the project was ultimately halted. It faced significant challenges, including low recall and poor accuracy for individuals with fair skin tones.
Figure 10. GHI forecasting customized DRGM
Figure 11. GHI forecasting Dataset#1 Evaluation Model
Figure 12. GHI forecasting Dataset#2 Evaluation Model
4.2. Hourly Global Horizontal Irradiance (GHI) Forecasting
@@
Context. Researchers in (9718423, ) investigated the use of LSTM models to forecast hourly Global Horizontal Irradiance (GHI) for a specific city, a critical attribute in determining the amount of solar energy that can be harvested from solar panel farms.
@@
Customizing the DRGM. First, the ML Problem Type-Based UML Activity Diagram was followed. Since this is a time-series regression problem, neither the balancedness of the data nor its redundancy was considered relevant. The amount of data KPI was set as recommended. Next, the Context-Based UML Activity Diagram was applied. As the data does not pertain to human subjects, directly impact human subjects, or contain sensitive information, the importance of data ethics requirements was set to ”None.” On the other hand, because the data is time-series in nature, the importance of Data Freshness and Data Management was set to ”High.” A temporal representativeness task was added to ensure seasonality is adequately captured. The resulting customized DRGM is shown in Fig. 10.
@@
Dataset Evaluation. The data used in (9718423, ) was obtained from the official website of K.A.CARE  [^2] https://rratlas.energy.gov.sa/, the organization responsible for installing and maintaining GHI monitoring stations across Saudi Arabia. This source is considered sustainable, ensures data freshness, and is regarded as authoritative. Incoming data was set to be handled manually. The importance of data ethics requirements was set to ”None” and assigned an initial value of $100$ during the evaluation to prevent negative impacts on the overall assessment. Some inconsistencies were identified but addressed during preprocessing. However, there was a 10% uncertainty in the measurement instruments, which affected the data’s accuracy. As the study in (9718423, ) focused on identifying the minimum data period required to achieve excellent performance (characterized by less than 10% mRMSE), the researchers experimented with various data volumes. They determined that a minimum of two years of data was necessary for representativeness. The dataset evaluation at these values is illustrated in Fig. 11.
However, the satisfaction of the Data Actor with this limited time frame is contingent on fully capturing temporal fluctuations. When replicating the experiment for another city, Najran, as shown in Fig. 12, the results did not hold. This was due to missing data for several months and the two years of training data not being representative of Najran’s overall climate, as they included atypical precipitation values.
5. Conclusion and Future work
The breakthroughs of ML in real-life applications such as medical diagnosis, banking, energy, and various other domains have highlighted the increasing need to address data requirements, as they are an essential building block of data-driven systems. Consequently, this work aimed to guide non-experts in ML in identifying the essential data requirements based on the ML problem they are attempting to solve.

### v2

@@ III-C Customization Mechanism of the DRGM
The customization mechanism is designed to adapt the initial DRGM to different ML problems and contexts. In addition, it assists requirements engineers, particularly those who are non-experts in ML, in selecting between datasets or evaluating a dataset to ensure that it satisfies the data requirements for a given ML problem, as illustrated in Fig. 4.
Figure 4: The Customization Mechanism
@@
Given that data requirements are highly dependent on the type of ML problem and are often context-specific, the customization mechanism categorizes data requirements goals into three sets. The first set includes goals that are of high importance across all ML problems, referred to as goal set #1. The second set comprises goals and subgoals that vary based on the ML type (e.g., Regression, Classification, Time Series). The third set pertains to data requirements that vary based on the collected data and the specific context. The contents of each set are detailed in Table II. Tasks can be added as illustrated in Fig 3.
TABLE II: The customization mechanism sets
@@
III-C1 Goals with Fixed Importance
@@
Examining the factors that determine set #1, Data Quantity is consistently assigned a ’High’ importance, as having more high-quality data generally leads to better results [47]. To satisfy the Data Quantity goal, both Data Availability and Data Accessibility must also be satisfied. Data Quality is equally critical; if Data Quality is not met, the acquired data will be ineffective, resulting in a ’garbage-in, garbage-out’ scenario, as noted by a data scientist in the study [47]. Similarly, Data Representativeness, Data Completeness, and Data Accuracy hold high importance across all contexts. That is, for Data Representativeness, a dataset must reflect the underlying population or environment it aims to model. Lack of representativeness leads to sampling bias, which causes poor generalization and fairness issues in downstream models [46]. Moreover, Data Completeness must be ensured to enable robust decision-making as missing values can lead to inaccurate or unstable models [31]. Noisy input or incorrect labels have been shown to reduce accuracy and affect model convergence [20], hence, Data Accuracy holds high importance as well. Although Data Consistency and Data Safety are also important, their significance is comparatively lower, and their importance is therefore set to ’Medium’. This is because they are often context-dependent [6]. That is, they are critical in regulated or adversarial settings but typically of medium importance in performance evaluation without such constraints [6].
III-C2 Goals with Different Importance Based on ML Problem Type
@@
Two elements related to data requirements vary in importance depending on the type of ML problem: Data Balancedness and Data Availability. Data Balancedness is highly important for classification problems but holds moderate importance for regression problems. For classification problems, techniques such as SMOTE, oversampling, and undersampling can be applied to address imbalanced data [11, 52]. For regression problems, data transformations to handle imbalance, often referred to as skewness, can be used [43, 22]. To set KPIs for Data Balancedness in classification problems, users should evaluate the balance as a percentage from the ratio (e.g., for binary classification, a 1:1 class ratio corresponds to 100% balance). For regression problems, the Shapiro-Wilks test is suggested to detect imbalances [35]. This test’s null hypothesis states that the data is drawn from a normally distributed population. If $P>0.05$, the null hypothesis is accepted, and the data is considered normally distributed. If the null hypothesis is rejected, the target variable is highly skewed [36]. The valuation values are determined based on the percentages or test results.
@@
For setting KPIs for Data Availability, several approaches exist to calculate the required amount of data based on the learning rate [19]. However, these approaches require an initial dataset and a pre-built ML algorithm. As such, they are primarily useful for determining the additional data needed during a pilot study for a selected dataset. When no initial dataset exists, there are no solid, theoretically proven rules for determining how much data an ML algorithm requires for training. Nonetheless, a heuristic known as the ’rule of 10,’ based on practitioners’ experiences, is frequently mentioned in the gray literature [24, 10, 37] and in a white paper for classification problems [40]. This rule suggests using ten instances per class or ten instances per predictor. Variations of this rule propose reducing the requirement to five instances or increasing it to 100. For this work, the KPI is set following this rule, incorporating its three variations which are classification, regression, and time series.
@@
In the case of image classification using deep learning, if a pre-trained model is not used, 1000 images per class are typically required as the minimum threshold [48]. Additional KPI values are derived from practitioners’ experiences and estimates. For forecasting with time-series data, particularly in seasonal data (e.g., weather or sales data), authors in [26] suggest using the number of seasons within a year plus five additional data points. However, real-world scenarios often involve randomness, requiring the model to utilize multiple ’seasons’ worth of data. Based on [50], the worst-case requirement is set to one times the number of data points in a season, the threshold to two times, and the maximum to ten times the number of data points in a season. For univariate, non-seasonal data using statistical models, the latest 50 observations are generally sufficient for short-term forecasts [34, 29]. Some practitioners from the gray literature suggest $40$ observations may suffice, while others recommend using up to $100$ for more accurate results. For multivariate time-series predictions, the ’rule of 10’ is also applied to refine the estimate. A detailed explanation of how to set the data size KPI for various ML problem types is provided in Table III, and the process is summarized in the MLProblemTypeBased UML activity diagram shown in Fig. 5.
TABLE III: KPI values for data size
@@
Classification – Image 500 * #classes 1000 * #classes 10,000 * #classes Number of images
@@
Time Series - Seasonal 1 2 10 Years (or the maximum seasonality period) worth of data
@@
Figure 5: ML Problem Type Based UML Activity Diagram
III-C3 Goals with Different Importance Related to Context
Many data requirements are context-dependent. For the Data Representativeness goal, the dataset should adequately capture variations across different geographic regions (spatial data) and time periods or seasons (seasonal-temporal data), to ensure the model can generalize across diverse conditions. If the prediction outcome directly impacts human subjects (e.g., in healthcare, criminal justice, or employment), it is critical to ensure Data Representativeness across targeted demographics. Lack of representativeness can lead to biased model predictions, resulting in discriminatory outcomes for underrepresented groups.
@@ III-C3 Goals with Different Importance Related to Context
Regarding Data Management, its importance is set to ’Low’ only if the model is built once and never updated. In such cases, Data Freshness is also set to ’Low’. However, a study by [13] found that one-third of models require updates at least monthly, and nearly one-quarter require daily tuning. This is especially relevant in fields such as marketing, stock market forecasting, and short-term weather prediction. For models requiring regular updates, Data Management is assigned ’High’ importance. Similarly, Data Freshness is set to ’High’ for models requiring frequent tuning, and the data must be acquired from sustainable sources that are ethically, legally, and practically viable for long-term use in ML without violating privacy, exhausting resources, or causing harm. For models with irregular updates, the importance of Data Freshness and sourcing sustainable data is set to ’Medium’.
@@
When considering Data Ethics, the content of the dataset must first be evaluated. If the data identifies human subjects, identifying information must be removed or protected to prevent discrimination and ensure the privacy. Additionally, user consent must be obtained as a task under Data Privacy. GDPR regulations impose stricter consent requirements for identifying information, such as a photograph of a person’s house. These rules apply when the subjects are European Union (EU) citizens or if the system will be used within the EU [1]. In other regions, compliance should align with the relevant local regulations.
@@
Next, we examine the Data Privacy goal. If the dataset contains sensitive information, such as health records, financial records, university records, or business records, the importance of both Data Privacy and Data Security is set to ’High’. If the data is private but not sensitive, their importance is set to ’Medium’. For public data, the importance of Data Privacy and Data Security is set to ’None’ as not all public data is directly usable. For private data, a release form must be signed to obtain access. For medical data, approval from an International Review Board (IRB) is required. Additionally, the importance of the goal Data Free from Discrimination is set to ’High’ if the ML model’s decisions have a significant impact on human lives, such as in job recruitment [15] or prison release [12]. In such cases, data should be obtained from authoritative sources, and sensitive demographic fields should either be removed or designated as protected to prevent bias in the system. Finally, the overall importance of Data Ethics is determined by the highest importance level among its subgoals. Before initiating the evaluation, an analyst should consult with both domain and legal experts to address any potential misrepresentation in the data and any legal concerns specific to the field. A summary of this process is illustrated in the ContextBased UML activity diagram, Fig. 6.
Figure 6: Context Based Activity Diagram
III-C4 Dataset Evaluation and Selection
@@
After customizing the DRGM based on the problem type and context, the dataset(s) are evaluated using the GRL evaluation strategy to select the dataset that best satisfies the Data Actor. We propose an evaluation scale ranging from 0 to 100. During the evaluation, all unrelated leaf elements should have their qualitative evaluation values set to ’satisfied,’ and their negative contribution links removed to avoid negatively impacting the overall evaluation.
@@
IV Illustrative Examples for Using DRGM
@@
This section shows the customization of the DRGM approach using the proposed customization mechanism. Since the data requirements depend on the ML prediction type and are mostly context-related, we choose two different illustrative examples. The first one, in the medical field, uses classification as the ML type. The second example uses the regression for time series weather forecasting.
Figure 7: Anemia Detection Customized DRGM
Figure 8: Anemia Detection Dataset#1 Evaluation Model
Figure 9: Anemia Detection Dataset#2 Evaluation Model
IV-A Non-invasive Diagnosis of Anemia System using Machine Learning
@@
Context. Researchers in [49] aimed to build an ML system to diagnose anemia using a video of a patient’s fingertip. To collect the training and testing dataset, three infrared LEDs and one white LED were positioned around the camera capturing the video. The LEDs were activated sequentially, and the average values of each of the three channels (RGB) from the video were calculated separately. These values were then used to train an ML model with the target classes being severely anemic, moderately anemic, mildly anemic, or healthy in a multi-class classification problem.
@@
Customizing the DRGM. First, we follow the ML problem type based UML activity diagram and as we are dealing with a classification problem we set the balancedness to ’High’, added the KPI as per the diagram, then added the related task. As for the KPIs for the data availability, we have $4$ classes, so we set the worst value to $20$, threshold value to $40$, and target value to $400$. We then follow the context-based UML activity diagram, and as the data relates to human subject Data Ethics requirements will be set to ’High’. No special legal requirements other than the IRB and the consent release forms are required. Different age groups, both female and male, and different skin colors should be represented. By completing this step, we now have a customized DRGM based on the ML type and based on the context, Fig. 8.
@@
Dataset Evaluation. The initial plan for the dataset involved collecting data in-house from 100 participants. Due to time constraints, the goal was to create a representative dataset and address any imbalance through sampling. Redundant data would be removed, and since the model was intended to be built only once, data would not be sourced from sustainable resources. Evaluating the planned dataset using the customized DRGM, shown in Fig. 8, indicated promising outcomes if the implementation were executed properly. However, as the implementation of the project described in [49] progressed, the total number of participants was reduced to $80$. The conventional participant selection process led to a dataset that lacked adequate representation of individuals with fair skin tones, male participants, and those who were moderately or severely anemic. After performing undersampling on the healthy class, the dataset remained highly imbalanced, with class ratios of 3:5:12:40. Consequently, the balancedness KPI was set to $20$. Certain inconsistencies in the dataset could not be resolved, as recordings were conducted in different rooms. Additionally, some accuracy issues arose due to the fading of lights towards the end of the data collection phase. The presence of multiple participants in the recording rooms further posed risks to data safety. The evaluation of the collected dataset, presented in Fig. 9, highlights why the project was ultimately halted. It faced significant challenges, including low recall and poor accuracy for individuals with fair skin tones.
Figure 10: GHI Forecasting Customized DRGM
Figure 11: GHI Forecasting Dataset#1 Evaluation Model
Figure 12: GHI Forecasting Dataset#2 Evaluation Model
IV-B Hourly Global Horizontal Irradiance (GHI) Forecasting
@@
Context. Researchers in [50] investigated the use of LSTM models to forecast hourly Global Horizontal Irradiance (GHI) for a specific city, a critical attribute in determining the amount of solar energy that can be harvested from solar panel farms.
@@
Customizing the DRGM. First, the ML Problem Type-Based UML Activity Diagram was followed. Since this is a time-series regression problem, neither the balancedness of the data nor its redundancy was considered relevant. The amount of data KPI was set as recommended. Next, the Context-Based UML Activity Diagram was applied. As the data does not pertain to human subjects, directly impact human subjects, or contain sensitive information, the importance of data ethics requirements was set to ’None’. On the other hand, because the data is time-series in nature, the importance of Data Freshness and Data Management was set to ’High’. A temporal representativeness task was added to ensure seasonality is adequately captured. The resulting customized DRGM is shown in Fig. 10.
@@
Dataset Evaluation. The data used in [50] was obtained from the official website of K.A.CARE  [^2] https://rratlas.energy.gov.sa/, the organization responsible for installing and maintaining GHI monitoring stations across Saudi Arabia. This source is considered sustainable, ensures data freshness, and is regarded as authoritative. Incoming data was set to be handled manually. The importance of data ethics requirements was set to ’None’ and assigned an initial value of $100$ during the evaluation to prevent negative impacts on the overall assessment. Some inconsistencies were identified but addressed during preprocessing. However, there was a 10% uncertainty in the measurement instruments, which affected the data’s accuracy. As the study in [50] focused on identifying the minimum data period required to achieve excellent performance (characterized by less than 10% mRMSE), the researchers experimented with various data volumes. They determined that a minimum of two years of data was necessary for representativeness. The dataset evaluation at these values is illustrated in Fig. 11. However, the satisfaction of the Data Actor with this limited time frame is contingent on fully capturing temporal fluctuations. When replicating the experiment for another city, Najran, as shown in Fig. 12, the results did not hold. This was due to missing data for several months and the two years of training data not being representative of Najran’s overall climate, as they included atypical precipitation values.
V Conclusion and Future work